• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 14:41:53 on October 26,1998

    Entry number 3733

    keyword=change to transverse pumping

    14:30 moved back to pol. He3 position after NMR measurement.
    Turned off longitudinal lasers, rotated field from 0 to 270 dgree.
    14:38-14:43 Turned on transverse lasers (4, 7 and 8).